Keira Mfg. is considering a rights offer. The company has determined that ex-rights price would be $71. The current price is $76 per share and there are 19 million shares outstanding. The rights offer would raise a total of $60 million. What is the subscription price?
Answer #1
Current market value = Current price * Current # of shares outstanding = 76 * 19
= $1,444 million
New market value = Current market value + Amount raised = 1444 + 60 = $1,504 million

Ex rights price = New market value / New # of shares outstanding
New # of shares outstanding = 1504 / 71 = 21.18 million

Number of shares issued in the rights offer = 21.18 - 19 = 2.18 million

Subscription price = Amount raised / Shares issued = 60 / 2.18 = $27.48 per share